# SeSE: A Structural Information-Guided Uncertainty Quantification Framework for Hallucination Detection in LLMs

### v2

@@ II-B Naive Structural Entropy
where $g_{\alpha}$ is the sum of weights of edges from the nodes in $\mathcal{T}_{\alpha}$ to those outside of $\mathcal{T}_{\alpha}$, $\mathcal{V}_{\alpha}=\sum_{v\in T_{\alpha}}d_{v}$ is the volume of $\mathcal{T}_{\alpha}$, and $\alpha^{-}$ is the parent node of $\alpha$.
II-C Predictive Uncertainty.
LLM generation can be viewed as a classification problem with extremely high dimension [19]. In classification, uncertainty is typically measured by predictive entropy (PE) [14]. The predictive entropy of the output random variable $Y$ with realization $y$ given $x$ is formally defined as $H(Y\mid x)=-\int p(y\mid x)\log(p(y\mid x))dy$, which becomes the following uncertainty score in natural language generation:
$U(x)=H({S}\mid x)=-\sum_{{s}}p({s}\mid x)\log(p({s}\mid x)).$ (4)
Here, the summation covers all possible generated sequences $s\in S$. High entropy indicates that many potential outputs have similar likelihoods, suggesting LLMs possess significant uncertainty about the correct answer and are highly likely to generate hallucinations. Conversely, low entropy implies a heavily concentrated output distribution, implying that LLMs probably possess relevant knowledge of correct answers.
Predictive uncertainty is occasionally characterized as total uncertainty, comprising both epistemic and aleatoric uncertainty [11]. Aleatoric uncertainty (data uncertainty) represents the irreducible uncertainty inherent in the data distributions, whereas epistemic uncertainty (model uncertainty) reflects the knowledge limitations of LLMs. Following most prior work, we focus on estimating total uncertainty, since both components jointly contribute to hallucinations. Although the distinction between these uncertainties would certainly enhance the performance of our proposed SeSE, we leave this extension to future work.
III Methodology
Figure 2: Overview of SeSE in sentence-length generation. For the sake of the example, we assume that the probabilities of each semantic cluster in (b) are equal, i.e., $p(c_{i}|\text{input})=p(c_{i}^{{}^{\prime}}|\text{input}),\forall c\in C$. Although the semantic structure of $\text{LLM}_{1}$ is more regular and its uncertainty should be lower (i.e., $\text{LLM}_{1}$ is quite certain that “urban greening” is a good answer), the semantic entropy fails to distinguish the differences and yields identical uncertainty scores due to their initially similar semantic spaces. In contrast, SeSE explicitly considers structural information via constructed semantic graphs, where concept “A” comprises two 3-level substructures “A.1” and “A.2”, thus correctly identifying that $\text{LLM}_{1}$ should be assigned lower uncertainty.
Figure 3: Overview of SeSE in long-form generation. We decompose the generated long-form response into atomic claims. SeSE quantifies the uncertainty of individual claims by modeling their random semantic interactions, and hallucination is indicated by a high SeSE value associated with that claim in the constructed bipartite response-claim graph.
This section provides an overview of our proposed SeSE, as illustrated in Figures 23. First, we formally define the semantic space and directed structural entropy in Section III-A, which form the theoretical foundation for SeSE. Then, we introduce the pipeline of SeSE for sentence-level generation ($16\pm 11$ tokens, mean $\pm$ s.d.) in Section III-B. Specifically, Section III-B1 specifies the sampling setup; Section III-B2 details the semantic graph construction algorithm; and Section III-B3 formally defines SeSE and introduces the proposed directed structural entropy minimization method. Finally, we elaborate on how SeSE operates at the claim-level in long-form generation ($142\pm 25$ tokens) in Section III-C.
III-A Semantic Space and Directed Structural Entropy
A semantic space models a set of interacting semantic entities and their dynamic relationships within a specific context. Note that each context defines a unique semantic space, i.e., identical semantic entities may exhibit different relationships across different contexts. Formally, such systems can be described using two equivalent mathematical frameworks: a graph-theoretic structure and an algebraic matrix model. We first define the graph model as follows:
Definition 4 (Graph Model of a Semantic Space).
A semantic space is formulated as a directed weighted graph $G_{dir}=(V,E,W)$, where:
  • $V$ is a finite set of vertices representing semantic entities.
  • $E\subseteq V\times V$ is a set of directed edges encoding interactions between vertices.
  • $W:E\to\mathbb{R}^{\geq 0}$ is a context-specific weight function such that $W(i,j)>0$ quantifies the interaction strength from semantic entities $i$ to entity $j$ within specific context. If $(i,j)\notin E$, then $W(i,j)=0$.
Weights can be instantiated using text embedding models, natural language inference models, or a weighted combination of the two, ensuring $W(i,j)\geq 0$ for all $i,j$.
The algebraic model enables the analysis of semantic uncertainty using linear algebra and spectral methods, whose formal definition is as follows:
Definition 5 (Algebraic Model of a Semantic Space).
A semantic space can be equivalently represented by a non-negative matrix $A\in\mathbb{R}_{\geq 0}^{|V|\times|V|}$, where each element $a_{i,j}=W(i,j)$ denotes the interaction from entity $i$ to $j$. The matrix $A$ thus fully preserves the structure of $G_{dir}$.
To address the limitations imposed by undirected constraints in the naive structural entropy [29, 32, 33], we define the high-dimensional structural entropy for directed graphs, enabling an accurate representation of directional semantic information in the semantic space. We first present the following theorem, forming the basis for the definition of directed structural entropy.
Theorem 1.
Given a directed graph $G_{dir}$ with non-negative edge weights and its corresponding matrix $A$, if $A$ is an irreducible stochastic matrix, then $G_{dir}$ possesses a unique stationary distribution $\pi^{\top}$ satisfying $\pi^{\top}A=\pi^{\top}$. This distribution corresponds to the unique eigenvector associated with the dominant eigenvalue 1 of the adjacency matrix $A$.
Proof of this theorem based on the Perron-Frobenius theorem is provided in Appendix VIII-A.
Therefore, to ensure the constructed directed semantic graph $G_{dir}$ possesses a unique stationary distribution $\pi$—a prerequisite for defining directed structural entropy—we define the adjusting operator. As presented in Algorithm 1, this operator serves two purposes: (i) ensuring $G_{dir}$ is irreducible by guaranteeing a directed path exists between any pair of vertices; and (ii) normalizing edge weights such that the weighted out-degree of each vertex sums to 1, thereby transforming $G_{dir}$ into a Markov chain conducive to a random walk. Specifically, we first identify the graph’s strongly connected components (SCCs) using Tarjan’s algorithm [34]. We then introduce directed edges with minimal weights to establish connectivity between these components. Finally, for each vertex $v\in V$, we normalize the weights of its outgoing edges by dividing each edge weight by the vertex’s weighted out-degree sum.
Input: A directed weighted graph $G_{dir}=(V,E,W)$ with non-negative edge weights
Output: A strongly connected directed graph $G_{dir}^{\prime}$ with normalized edge weights
1
1ex// SCCs decomposition
2 $SCCs\leftarrow\text{Tarjan}(G_{dir})$ [34], $G_{SCC}=(V_{SCC},E_{SCC})$
1ex// make $G_{SCC}$ strongly connected
3 $sources\leftarrow\text{SCCs with no incoming edges in }G_{SCC}$
4 $sinks\leftarrow\text{SCCs with no outgoing edges in }G_{SCC}$
5 if $|V_{SCC}|>1$ then
6 foreach sink SCC $S_{sink}$ do
7    foreach source SCC $S_{source}$ do
8       if there’s no path from $S_{sink}$ to $S_{source}$ then
9            Add a minimum-weight edge from $S_{\text{sink}}$ to $S_{\text{source}}$
10
11
12
13
1ex// out-degree normalization
14 foreach edge $(v_{i},v_{j})\in E$ do
15 $W(v_{i},v_{j})\leftarrow W(v_{i},v_{j})/\sum_{(v_{i},v_{k})\in E}W(v_{i},v_{k})$
return $G_{dir}^{\prime}\leftarrow G_{dir}$
Algorithm 1 Adjusting Operator
For the adjusted semantic graph $G_{dir}^{\prime}=(V,E^{\prime},W^{\prime})$, we compute the stationary distribution $\pi$ over all vertices and define the one-dimensional directed structural entropy as follows:
$H^{1}(G_{dir}^{\prime})=-\sum_{v\in V}\pi(v)\cdot\log_{2}\pi(v),$ (5)
where $\pi(v)$ denotes the stationary probability of vertex $v$ in $G_{dir}^{\prime}$. Using the stationary distribution $\pi$, we redefine $g_{\alpha}$ and $V_{\alpha}$ for each non-root node $\alpha$ in the encoding tree $\mathcal{T}_{dir}$ of $G_{dir}^{\prime}$, and its assigned entropy as follows:
$\mathcal{V}_{\alpha}=\sum_{v_{i}\in V}\sum_{v_{j}\in V_{\alpha}}\left[\pi\left(v_{i}\right)\cdot W^{\prime}\left(v_{i},v_{j}\right)\right],$ (6)
$g_{\alpha}=\sum_{v_{i}\notin V_{\alpha}}\sum_{v_{j}\in V_{\alpha}}\left[\pi\left(v_{i}\right)\cdot W^{\prime}\left(v_{i},v_{j}\right)\right],$ (7)
$H^{\mathcal{T}_{dir}}\left(G_{dir}^{\prime};\alpha\right)=-\frac{g_{\alpha}}{vol\left(G_{dir}^{\prime}\right)}\cdot\log_{2}\frac{\mathcal{V}_{\alpha}}{\mathcal{V}_{\alpha^{-}}},$ (8)
where the volume of $G_{dir}^{\prime}$ is redefined as the sum of in-degrees and out-degrees of all vertices:
$vol\left(G_{dir}^{\prime}\right)=\sum_{v\in V}\left(d_{v}^{+}+d_{v}^{-}\right),$ (9)
where $d_{v}^{+}$ and $d_{v}^{-}$ denote the weighted out-degree and in-degree of vertex $v$, respectively. Thus, the directed structural entropy of $G_{dir}^{\prime}$ is redefined as follows:
$H^{\mathcal{T}_{dir}}\left(G_{dir}^{\prime}\right)=\sum_{\alpha\in\mathcal{T}_{dir},\alpha\neq\lambda}H^{\mathcal{T}_{dir}}\left(G_{dir}^{\prime};\alpha\right),$ (10)
III-B SeSE for Sentence-length Generation
While semantic entropy aims to estimate the LLM’s uncertainty in the semantic space, it fails to capture the intrinsic semantic structure. Therefore, SeSE extends the simple Shannon entropy of semantic clusters to the $K$-dimensional structural entropy of directed semantic graphs, enabling more precise uncertainty estimation. At a high level, this involves three steps: response sampling, semantic graph construction and hierarchical abstraction.
III-B1 Response Sampling
Unlike white-box UQ approaches that require access to the internal token probabilities of LLMs, SeSE only needs a set of sampled answers as input, thus having broader applicability. Given the context $x$ as input to $\mathcal{M}$, we sample two types of responses from $\mathcal{M}$: a single greedy-decoded answer $r_{T=0}$ and a set of $N$ stochastically sampled answers $\mathcal{R}(\cdot\mid x)=\left\{r_{T=1}^{1},\ldots,r_{T=1}^{N}\right\}$. Responses in $\mathcal{R}$ are generated at a temperature of $T=1$ using nucleus (top-$P$) sampling [35] and top-$K$ sampling [36], with parameters $P=0.95,K=20$. The greedy decoding response $r_{T=0}$, obtained by setting $T=0,P=1,K=20$, serves as the model’s most confident answer and is used for accuracy evaluation. The temperature parameter $T$ controls the stochasticity of the LLM’s output by scaling the logits over the next token. A lower temperature value amplifies the probability of initial high-probability tokens. $T=1$ preserves the model’s original output distribution, whereas $T=0$ results in deterministic decoding by always selecting the highest probability token. We default to setting the sample size $N=10$ and the relevant hyperparameter analysis is presented in Section V-D1. The prompt template can be found in Appendix VIII-F.
III-B2 Semantic Graph Construction
Input: Context $x$, the sampled response set $\mathcal{R}(\cdot\mid x)=\left\{r_{T=1}^{1},\ldots,r_{T=1}^{N}\right\}$, NLI model
Output: An irreducible, directed sparsified semantic graph $G^{*}_{dir}=(V,E,W)$ with normalized edge weights
// pairwise entailment measurement
1 Initialize a weight matrix $A\in\mathbb{R}^{N\times N}$ to 0
2 foreach ordered pair $(r^{i},r^{j})\in\mathcal{R}\times\mathcal{R}$ do
3 if $i\neq j$ then
4      compute $P_{NLI}(r^{i}\rightarrow r^{j}|x)$ via Eq. 11
5      set $A_{i,j}$ via Eq. 12
6   end if
7
8 end foreach
// initialize semantic graph
9 $G_{dir}\leftarrow A$
// graph sparsification
10 for $k=1$ to N1N-1 do
11   construct $G_{k}$ as the kkNN graph from $G_{dir}$
12 $G_{k}\leftarrow\text{Adjusting}(G_{k})$ // apply Algorithm 1
13   calculate $H^{1}(G_{k})$ via Eq. 5
14 end for
@@ 12 $G_{k}\leftarrow\text{Adjusting}(G_{k})$ // apply Algorithm 1
return $G^{*}_{dir}\leftarrow G_{k^{*}}$
Algorithm 2 Proposed AS-DSG
Prior studies have generally constructed undirected, fully connected semantic graphs, which neglect inherently directional semantic information and often introduce numerous redundant connections. To capture directional relationships between responses and eliminate interference from irrelevant connections, we propose an adaptively sparsified directed semantic graph construction algorithm (AS-DSG). The main input to AS-DSG is the response set $\mathcal{R}$ sampled in Section III-B1, and its output is an irreducible (i.e., strongly connected), directed sparsified semantic graph $G^{*}_{dir}=(V,E,W)$ with normalized edge weights. As presented in Algorithm 2, AS-DSG proceeds in two stages: pairwise entailment measurement and graph sparsification.
Pairwise entailment measurement: First, we perform pairwise semantic entailment measurement to obtain the mathematical representation of the semantic space, which facilitates the subsequent analysis of semantic uncertainty (lines 2-7 in Algorithm 2). Most existing text embedding models exhibit limitations in measuring contextual semantic similarity. Given a context $x$ and a response $r$, they typically generate an embedding for the whole concatenated sequence $E(x\oplus r)$ instead of a context-conditioned embedding $E(r|x)$. In contrast, natural language inference (NLI) models have been demonstrated to effectively analyze contextually semantic relationships between texts [16].
To preserve semantic directional information, we employ the NLI model DeBERTa-v3-large-mnli [^2] https://huggingface.co/khalidalt/DeBERTa-v3-large-mnli to perform pairwise entailment measurements on the sampled response set $\mathcal{R}$. For each ordered pair $(r^{i},r^{j})\in\mathcal{R}\times\mathcal{R}$ with $r^{i}\neq r^{j}$, we construct a premise–hypothesis pair $(x\oplus r^{i},x\oplus r^{j})$. Given each pair as input, the NLI model outputs corresponding logits for three classes: entailment, neutral, and contradiction, capturing the directional semantic relationships from $x\oplus r^{i}$ (premise) to $x\oplus r^{j}$ (hypothesis). We then apply the softmax function $\sigma(\cdot)$ to these logits to obtain a probability distribution $P_{NLI}(r^{i}\rightarrow r^{j}|x)$, which is denoted as a three-element tuple:
$\left[{p_{e}}(r^{i}\rightarrow r^{j}|x),{p_{n}}(r^{i}\rightarrow r^{j}|x),{p_{c}}(r^{i}\rightarrow r^{j}|x)\right]$
$=\sigma\left(NLI(x\oplus r^{i},x\oplus r^{j})\right).$ (11)
Here, ${p_{e}}(r^{i}\rightarrow r^{j}|x),{p_{n}}(r^{i}\rightarrow r^{j}|x),{p_{c}}(r^{i}\rightarrow r^{j}|x)$ denote the semantically equivalent, irrelevant, or contradictory probabilities from $r^{i}$ to $r^{j}$ conditioned on the context $x$, respectively.
To this end, an asymmetric adjacency matrix $A$ is derived for constructing the directed semantic graph $G_{dir}=(V,E,W)$. The $V=\mathcal{R}$ denotes the set of sampled responses and $E$ is the set of directed edges with weights defined by the context-specific function $W$. The edge weight from node $i$ to node $j$ is defined as:
$W(r^{i},r^{j})=A_{i,j}=\omega\cdot P_{NLI}(r^{i}\rightarrow r^{j}|x),$ (12)
where $\omega=(1,\dfrac{1}{2},0)^{\top}$ is a weight vector. This value $A_{i,j}$ represents the probability of a directed semantic transformation from response $r^{i}$ to $r^{j}$. Unless the $r^{i}$ and $r^{j}$ are completely identical, $A_{i,j}$ is not necessarily equal to $A_{j,i}$. This asymmetry distinguishes $G_{dir}$ from traditional undirected semantic graphs, which often impose symmetry by averaging bidirectional edge weights (e.g., $\dfrac{A_{i,j}+A_{j,i}}{2}$). Such averaging can obscure the directionality of semantic relations and result in information loss.
Graph sparsification: While constructing a semantic entailment graph to represent all inter-node relationships is reasonable, the resulting initial graph is fully connected (i.e., all nodes belong to a single class). This configuration introduces numerous insignificant connections that obscure the underlying semantic structure. To address this, we propose an adaptive sparsification mechanism (lines 12–16 in Algorithm 2). Based on structural information theory [31], we sparsify the graph $G_{dir}$ into a kkNN graph $G_{k}$ by minimizing its one-dimensional structural entropy. Notably, $k$ is determined adaptively based on structural entropy, requiring no manual specification. Specifically, the sparsification procedure involves treating each sampled response $r\in\mathcal{R}$ as a central vertex and retaining only the highest $k$ outgoing edge weights to construct a candidate kkNN graph $G_{k}$ and compute corresponding one-dimensional structural entropy $H^{1}(G_{k})$. (The adjusting operator ensures the existence and uniqueness of the stationary distribution $\pi$ of $G_{k}$). We evaluate $H^{1}(G_{k})$ across a reasonable range of $k$ values, select optimal parameter $k^{*}$ that satisfies the following inequality:
$H^{1}(G_{k^{*}-1})\geq H^{1}(G_{k^{*}})\leq H^{1}(G_{k^{*}+1}),$ (13)
and output $G_{k^{*}}$ as the sparsified semantic graph $G^{*}_{dir}$, which preserves the most useful structural information from the initial complete graph.
III-B3 Hierarchical Abstraction
Input: the sparsified directed semantic graph $G^{*}_{dir}$, $K>1$
Output: the $K$-dimensional optimal encoding tree $\mathcal{T}^{*}$
1 Initialize a one-dimensional encoding tree $\mathcal{T}$
2 while tree height $h<K$ do
3 foreach sibling nodes $\alpha,\beta\in\mathcal{T}$ do
@@ 1 Initialize a one-dimensional encoding tree $\mathcal{T}$
5   end foreach
6 if $\Delta SeSE_{\alpha,\beta}^{op_{mer}}>0$ then
7    $\mathcal{T}\leftarrow op_{mer}(\mathcal{T},\alpha^{*},\beta^{*})$, continue
8   end if
9 foreach sibling nodes $\alpha,\beta\in\mathcal{T}$ do

### v3

@@ 2 Preliminaries
where $g_{\alpha}$ is the total weight of edges from nodes in $\mathcal{T}_{\alpha}$ to nodes outside $\mathcal{T}_{\alpha}$; $\mathcal{V}_{\alpha}$ denotes the volume of the sub-graph induced by $\mathcal{T}_{\alpha}$, i.e., the sum of degrees of its constituent vertices; and $\alpha^{-}$ denotes the parent node of $\alpha$.
3 Semantic Structural Entropy
Figure 1: Illustration of SeSE in short-form generation. Superficially, both semantic spaces in (2) appear similar, as they each contain three semantic substructures. However, the hierarchical abstraction of $\text{LLM}_{1}$ is notably more regular, featuring a hierarchically nested structure where the high-level substructure A aggregates two finer-grained substructures, enabling efficient compression. In contrast, the hierarchical abstraction of $\text{LLM}_{2}$ is more disordered and resists compression. This implies $\text{LLM}_{1}$ exhibits lower uncertainty than $\text{LLM}_{2}$, i.e., $\text{LLM}_{1}$ is more certain that "greening" is an excellent solution. SeSE captures the inherent semantic structure by constructing the optimal hierarchical abstraction, which is formalized by an encoding tree with minimal structural entropy, thereby correctly assigning $\text{LLM}_{1}$ a lower uncertainty score.
3.1 SeSE for Short-form Generation
As shown in Fig. 1, SeSE comprises three phases: response sampling, semantic graph construction and hierarchical abstraction. In Step 3, we construct the optimal semantic hierarchical abstraction based on the structural entropy minimization principle Li and Pan [2016]. The structural entropy of the resulting encoding tree quantifies the inherent uncertainty of the semantic space after optimal compression. Semantic spaces with clear regularities exhibit lower entropy, corresponding to lower uncertainty, while disordered spaces resist compression and yield higher entropy.
Step 1. Response Sampling
SeSE is a sampling-based approach that requires only a set of black-box responses as input. Given the context $x$ as input to an LLM $\mathcal{M}$, we sample two types of responses from $\mathcal{M}$: (1) a single greedy-decoded answer $r_{T=0}$ generated at temperature $T=0$, which serves as a proxy for the model’s most confident output; and (2) a set of $N$ stochastic samples $R=\left\{r^{1},\ldots,r^{N}\right\}$ for semantic space modeling.
Step 2. Semantic Graph Construction
Natural language inference (NLI) models have proven to be effective in analyzing contextual semantic relationships between texts Farquhar et al. [2024]. We employ the NLI model DeBERTa-v3-large-mnli He et al. [2021] to measure pairwise entailment within the response set $R$. For each ordered pair $(r^{i},r^{j})\in R\times R$ with $i\neq j$, we construct a premise-hypothesis pair $(x\oplus r^{i},x\oplus r^{j})$ and feed it into the NLI model. We then apply the softmax function $\sigma(\cdot)$ to the predicted logits to obtain a probability distribution:
$p_{e},p_{n},p_{c}]=\sigma\left(\text{NLI}(x\oplus r^{i},x\oplus r^{j})\right),$ (6)
where $p_{e}$, $p_{n}$, and $p_{c}$ denote the probabilities of entailment, neutrality, and contradiction from the premise to the hypothesis, respectively, conditioned on $x$. Formally, we model the semantic space using a directed weighted graph $G=(V,E,W)$ where the vertex set is $V=R$ and $E$ is the set of directed edges. The weight $W(v_{i},v_{j})$ is defined as
$W(v_{i},v_{j})=p_{e}+\frac{1}{2}\cdot p_{n}+0\cdot p_{c}=p_{e}+\frac{1}{2}\cdot p_{n}.$ (7)
To make $G$ suitable for a random-walk process, we normalize the weights of all outgoing edges by dividing each weight by the vertex’s weighted out-degree sum, obtaining the transition matrix $P$:
$P(v_{i},v_{j})=\frac{W(v_{i},v_{j})}{\sum_{(v_{i},v_{k})\in E}W(v_{i},v_{k})}.$ (8)
Step 3. Hierarchical Abstraction
Input: the directed semantic graph $G$, $K>1$
Output: the $K$-dimensional optimal encoding tree $\mathcal{T}^{*}$
Initialize a one-dimensional encoding tree $\mathcal{T}$ for $G$
while tree height $h<K$ do
 foreach sibling nodes $\alpha,\beta\in\mathcal{T}$ do
@@ 3.1 SeSE for Short-form Generation
   end foreach
 if $\Delta_{\alpha^{*},\beta^{*}}^{op_{mer}}>0$ then
    $\mathcal{T}\leftarrow op_{mer}(\mathcal{T},\alpha^{*},\beta^{*})$, continue
   end if
 foreach sibling nodes $\alpha,\beta\in\mathcal{T}$ do
@@
   end foreach
 if $\Delta_{\alpha^{*},\beta^{*}}^{op_{com}}>0$ then
    $\mathcal{T}\leftarrow op_{com}(\mathcal{T},\alpha^{*},\beta^{*})$, continue
   end if
  break
end while
return $\mathcal{T}^{*}\leftarrow\mathcal{T}$
Algorithm 1 Hierarchical Abstraction Construction
Since $P$ is an irreducible row-stochastic matrix with non-negative entries, there exists a unique stationary distribution $\pi$ satisfying $\pi P=\pi$ Norris [1998]. We then define and optimize structural entropy for directed graphs based on its stationary distribution, enabling accurate representation of key transition dynamics in semantic spaces. The one-dimensional directed structural entropy of $G$ is define as follows:
$H^{1}(G)=-\sum_{v\in V}\pi(v)\cdot\log_{2}\pi(v),$ (9)
where $\pi(v)$ denotes the stationary probability of vertex $v$. We then refine the definitions of $g_{\alpha}$ and $\mathcal{V}_{\alpha}$ for each non-root node $\alpha$ in the encoding tree $\mathcal{T}$. Following Eq. 5, the entropy assigned to $\alpha$ is defined as:
$\mathcal{V}_{\alpha}=\sum_{v_{i}\in V}\sum_{v_{j}\in V_{\alpha}}\pi(v_{i})P(v_{i},v_{j}),$ (10)
$g_{\alpha}=\sum_{v_{i}\in V\setminus V_{\alpha}}\sum_{v_{j}\in V_{\alpha}}\pi(v_{i})P(v_{i},v_{j}),$ (11)
$H^{\mathcal{T}}(G;\alpha)=-\frac{g_{\alpha}}{\text{vol}(G)}\log_{2}\frac{\mathcal{V}_{\alpha}}{\mathcal{V}_{\alpha^{-}}},$ (12)
where $\text{vol}\left(G\right)=\sum_{v\in V}\left(d_{v}^{+}+d_{v}^{-}\right)$ is the volume of $G$, with $d_{v}^{+}$ and $d_{v}^{-}$ denoting the weighted out-degree and in-degree of vertex $v$, respectively.
The optimal hierarchical abstraction is determined by identifying an encoding tree $\mathcal{T}^{*}$ with minimal structural entropy Li [2024]. To find $\mathcal{T}^{\star}$, we design an efficient structural entropy minimization algorithm (Algorithm 1) using the merging ($op_{mer}$) and combining ($op_{com}$) operators introduced by Li [2024]. Let $\mathcal{T}_{\alpha,\beta}$ be the encoding tree obtained after executing the merging or combining operator on sibling nodes $\alpha$ and $\beta$ that share a common parent. We define the resulting entropy variation as
$\Delta_{\alpha,\beta}^{op}=H^{\mathcal{T}}(G)-H^{\mathcal{T}_{\alpha,\beta}}(G).$ (13)
Specifically, we first initialize a one-dimensional encoding tree $\mathcal{T}$ : (1) the root node $\lambda$ represents the entire semantic space, i.e., $\mathcal{T}_{\lambda}=R$; (2) for each $r\in R$, we generate a leaf node $\gamma$ with $\mathcal{T}_{\gamma}=\left\{r\right\}$ and assign it as a child node of $\lambda$, i.e., $\gamma^{-}=\lambda$. Subsequently, we iteratively optimize the encoding tree $\mathcal{T}$ to $K$ layers. In each iteration, we traverse all sibling node pairs in $\mathcal{T}$ and greedily apply $op_{mer}$ or $op_{com}$, selecting the operation that maximizes the decrease of structural entropy while ensuring that the tree height remains below $K$. The iterative procedure terminates when no sibling pair satisfies $\Delta_{\alpha,\beta}^{op}>0$ or the tree height reaches $K$, at which point we output $\mathcal{T}^{*}$. A detailed illustration of the operators and entropy variations is provided in Appendix C. Formally, we define SeSE as the total entropy of the optimal $K$-dimensional encoding tree $\mathcal{T}^{*}$:
$\mathcal{T}^{\star}=\underset{\forall\mathcal{T}:\operatorname{height}(\mathcal{T})\leq K}{\arg\min}\left(H^{\mathcal{T}}(G)\right),$ (14)
$Se$ $SE(G)=\sum_{\alpha\in{\mathcal{T}^{\star}},\alpha\neq\lambda}H^{\mathcal{T}^{\star}}(G;\alpha).$ (15)
SeSE Generalizes Semantic Entropy [Farquhar et al., 2024].
The following theorem shows that SeSE can recover SE for any semantic clustering.
Theorem 1 (SeSE Generalizes SE).
For any semantic clustering, there exists a semantic graph such that the one-dimensional structural entropy of the graph is equal to the semantic entropy (computed as in Eq. 2).
Proof Sketch. When $K=1$, SeSE reduces to the Shannon entropy of the stationary distribution of the graph (Eq. 9). Given any semantic clustering, we can construct a cluster-level semantic graph whose stationary distribution matches this cluster distribution. SeSE with $K=1$ of this constructed graph equals the SE. Appendix A provides the detailed proof.
3.2 SeSE for Long-form Generation
LLMs often output long-form paragraphs comprising multiple claims Min et al. [2023], which are the smallest semantically distinct information units. In long-form generation, we therefore assess uncertainty at the finer claim level rather than simply assigning a single uncertainty score to an entire response. Given a context $x$, a set of stochastic sampled responses $R$, and a set of claims $C$ extracted from the greedy response $r_{T=0}$, we construct a claim-response bipartite graph $G_{cr}=((R,C),E)$ where an edge $(r,c)\in E$ indicates that response $r$ semantically entails claim $c$.
By minimizing the $K$-dimensional structural entropy of $G_{cr}$ with Algorithm 1, we obtain its optimal encoding tree $\mathcal{T}_{cr}^{*}$, which captures the inherent hierarchical community structure over $R\cup C$. For any claim $c\in C$, the uncertainty of reaching $c$ is determined by the cumulative entropy of all non-root nodes $\alpha$ encountered along the path from the root $\lambda$ to the leaf $\gamma$ with $V_{\gamma}=\{c\}$. Accordingly, we define the SeSE of each claim $c$ as its uncertainty of engaging in random interactions within $G_{cr}$:
$\text{SeSE}\left(G_{cr};c\right)=-\sum_{\alpha\in\mathcal{P}(\lambda\to\gamma)\setminus\{\lambda\}}\frac{g_{\alpha}}{\operatorname{vol}(G_{cr})}\log_{2}\frac{\mathcal{V}_{\alpha}}{\mathcal{V}_{\alpha^{-}}},$ (16)
where $\mathcal{P}(\lambda\to\gamma)$ denotes the path from the root $\lambda$ to the leaf $\gamma$ representing $c$. Claims with lower SeSE reside in densely connected core regions, reflecting consistent support across sampled responses and lower uncertainty. Conversely, claims with higher SeSE lie in peripheral or sparsely connected regions, indicating a higher uncertainty. Implementation details are provided in Appendix B.
4 Experimental
4.1 Experimental Setup
Datasets and LLMs
For short-form experiments, we employ five representative free-form QA datasets spanning diverse domains of natural language generation: BioASQ Krithara et al. [2023] (biomedical sciences), SVAMP Patel et al. [2021] (mathematical word problems), TriviaQA Joshi et al. [2017] (trivia knowledge), NQ-Open Kwiatkowski et al. [2019] (open-domain natural questions), and SQuAD_V2 Rajpurkar [2018] (reading comprehension). The evaluation is conducted across the Llama-3.1-Instruct series (8B and 70B parameters) Meta [2024] and the Qwen-3-Instruct series (4B and 30B parameters) Qwen [2025]. Regarding long-form generation, we perform evaluations on two challenging datasets featuring real-world entities from Wikipedia: FActScore Min et al. [2023] and PopQA Mallen et al. [2023]. For these benchmarks, we utilize DeepSeek-V3.1 DeepSeek-AI [2025] and Gemini-3-Flash Google DeepMind [2025]. Further dataset details are provided in Appendix F.
Baselines
We include a range of widely used baselines for comparison. In short-form experiments, we evaluate seven representative UQ methods: (1) P(True) Kadavath et al. [2022] uses few-shot prompts to guide LLMs to estimate the probability that their most confident answer is true. (2) Embedding Regression (ER) Farquhar et al. [2024] is a strong supervised baseline that trains a classifier on the final hidden states to predict correctness. (3) SelfCheckGPT (SC) Manakul et al. [2023] is a representative self-consistency method. (4) Length-normalized Predictive Entropy (LN-PE) Malinin and Gales [2021] computes length-normalized token-level log-probabilities. (5) Semantic Entropy (SE) Farquhar et al. [2024] estimates the Shannon entropy over semantic clusters. (6) Kernel Language Entropy (KLE) Nikitin et al. [2024] is a generalization of SE using semantic kernels. (7) Semantic Graph Density (SGD) Li et al. [2025b] quantifies semantic consistency via semantic graph density. In long-form experiments, in addition to adapting DSE, SC, and P(True) for long-form generation, we further incorporate two Verbalized Uncertainty methods Mohri and Hashimoto [2024], including Post-hoc Verbalized Uncertainty (PH-VU) and In-line Verbalized Uncertainty (IL-VU). For further details, refer to Appendix G.
Evaluation Metrics
Following previous work Farquhar et al. [2024], we assess two primary metrics: Area Under the Receiver Operating Characteristic curve (AUROC) and Area Under the Rejection Accuracy Curve (AURAC) Nadeem et al. [2010]. AUROC evaluates how well the uncertainty scores distinguish between correct and incorrect answers and ranges from 0 to 1, where 1 denotes a perfect classifier and 0.5 indicates random classification. AURAC quantifies the potential accuracy improvement users may experience when employing different UQ metrics to exclude high-uncertainty queries. The X% rejection accuracy represents model performance on the subset of questions retained after filtering out the top X% high-uncertainty queries, and AURAC provides a comprehensive assessment of accuracy across multiple rejection thresholds.
Implementation Details
For KLE and SGD, we employ the best variants $\text{KLE}_{\text{HEAT}}$ and $\text{SGD}_{\text{s+P}}$, and use DeBERTa-v3-large-mnli as the NLI model, which is identical to SeSE. To ensure fairness, for SC, SE, DSE and SeSE, we use identical GPT-5-mini for entailment prediction. In all experiments, stochastic responses $R$ are generated at a temperature of $T=1$ with a size of $N=10$ using nucleus sampling ($P=0.95$) Holtzman et al. [2020] and top-$K$ sampling ($K=20$) Fan et al. [2018]. The greedy-decoded response $r_{T=0}$, obtained at $T=0,P=1,K=20$, serves as the most confident answer and is used for accuracy evaluation. We use GPT-5-mini to automatically evaluate the correctness of $r_{T=0}$ by comparing it with the reference answer. We validate this automated evaluation against human judgment, and relevant results are shown in Appendix D.5.
4.2 Main Results